OpenStudy (anonymous):

Find the solutions to the nonlinear equations y=4x x^2+y^2=68

OpenStudy (anonymous):

y=4x y^2=16x^2 x^2+16x^2=68 17x^2=68 x^2=68/17 x^2=4 x=+-2